5 pencils and 7 pens together cost Rs.195. While 7 pencils and 5 pens together cost Rs. 153. Find the cost of each pencil and the pen.
1
Expert's answer
2013-07-09T09:41:39-0400
Let x is the cost of each pencil, and y is the cost of each pen.
Then 5x + 7y = 195 and 7x + 5y = 153.
Solve this system of equations.
35x + 49y = 1365
35x + 25y = 765

24y = 600
y = 25

5x + 7*25 = 195
5x = 20
x = 4;

Answer: pencil costs Rs. 4, pen costs Rs. 25.